# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v3

@@ III.7 Analysis Details
Figure 3: $\hat{q}_{2}^{2}$ distributions. Shown are the single $\hat{q}_{2}^{2}$ and pair $\hat{q}_{2,{\rm pair}}^{2}$ distributions of POIs from avfd with $n_{5}/s=0$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I (lower center), and Toy Model II (lower right). Shown in all panels are for the 30–40% centrality of Au+Au collisions, as examples. The $q_{2}^{2}$ (used in the ESE analysis) is calculated from particles in $|\eta|<0.3$, and the pair $q_{2,{\rm pair}}^{2}$ (used in the ESS analysis) is calculated from particles within $|\eta|<1$, both with $20.2<p_{\rm T}<$ 2 GeV/$c$.
POIs are assumed to be all charged hadrons, defined to be $\pi^{\pm}$, $K^{\pm}$, $p$ and $\bar{p}$ in our study, and are within the transverse momentum range of $20.2<p_{\rm T}<$ 2 GeV/$c$ and a certain central pseudorapidity ($\eta$) range, typical of midrapidity detectors such as the STAR experiment [68].
@@
  • east subevent: $10.3<\eta<$ 1 (or $20.3<\eta<$ 2).
The POIs are taken from either the west subevent or the east subevent, but not from both. The event selection quantity $\hat{q}_{2}^{2}$ is computed by the two-particle cumulant method in Eqs. 5,6, using particles from the center subevent, $|\eta|<0.3$ (or from a forward/backward region $43<|\eta|<$ 4).
@@
Figure 4: Average $v_{2}^{2}\{2\}$. Shown are the two-particle cumulant $v_{2}^{2}\{2\}$ of POIs and $v_{2,{\rm pair}}^{2}\{2\}$ of particle pairs of POIs as functions of centrality from avfd for $n_{5}/s=0$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I (lower center), and Toy Model II (lower right). POIs are from acceptances $|\eta|<1$ (except for epos4, $|\eta|<2$) and $20.2<p_{\rm T}<$ 2 GeV/$c$. The large departure of $v_{2,{\rm pair}}^{2}\{2\}$ from $v_{2}^{2}\{2\}$, both calculated by the cumulant method (Eq. 8), is largely due to self-correlations where two pairs share a common particle. The issue is severe in peripheral collisions except for hydjet++ where the average multiplicity is significantly larger than the rest models; for reference, for 30–40% centrality of Au+Au, the average multiplicities within $|\eta|<0.5$ are 151, 169, 88, and 255 for avfd, ampt, epos4, and hydjet++, respectively.
@@
The $\hat{q}_{2}^{2}$ variable calculated from single particles has been widely used. The $\hat{q}_{2,{\rm pair}}^{2}$ variable calculated from particle pairs is relatively new [36]. As seen from Eq. 7, the average $\left\langle q_{2}^{2}\right\rangle$ is related to the two-particle cumulant $v_{2}\{2\}$. To gain insights, we calculate by Eq. 8 the $v_{2}\{2\}$ of single POIs and the $v_{2,{\rm pair}}\{2\}$ of particle pairs of POIs; In the latter case, the $\phi_{1}$ and $\phi_{2}$ in Eq. 8 are the azimuthal angles of two pairs, each formed by two POIs. The $v_{2,{\rm pair}}\{2\}$ consists of a component from pairing of “random” (hydrodynamic) particle pairs or a hydrodynamic pair with a resonance-decay pair, and a component from two-resonance $v_{2}\{2\}$. The first component reflects, in a non-trivial way, the single particle $v_{2}$ [69]. The last component is simply resonance (background contributing source) $v_{2}\{2\}$. Both $v_{2}\{2\}$ and $v_{2,{\rm pair}}\{2\}$ contain nonflow contamination, for example, from resonance decays (in the case of $v_{2,{\rm pair}}\{2\}$, the two pairs both contain daughters from the same resonance decay). Figure 4 shows $v_{2}\{2\}$ (from 2 POIs) and $v_{2,{\rm pair}}\{2\}$ (from 2 pairs of POIs) as functions of centrality in different simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200\text{ GeV}$. They roughly agree at high multiplicity (hydjet++, or central collisions of other models), and $v_{2,{\rm pair}}\{2\}$ becomes much larger than $v_{2}\{2\}$ at low multiplicity (peripheral collisions of models except for hydjet++). This discrepancy comes from the self-correlations where the two pairs share one common POI. To get a feeling of this effect, consider a simplified scenario with $\phi_{{\rm pair}}=(\phi_{1}+\phi_{2})/2$. For two pairs with indices $(1,3)$ and $(2,3)$ where particle 3 is shared, their correlation is $\cos 2[(\phi_{1}+\phi_{3})/2-(\phi_{2}+\phi_{3})/2]=\cos(\phi_{1}-\phi_{2})$, and the fraction of this case in the $v_{2,{\rm pair}}^{2}\{2\}$ calculation is the triplet multiplicity over the quadruplet one, which is $\sim 1/N$. Therefore, besides the common flow component, $v_{2,{\rm pair}}\{2\}$ has an extra part $\sim\langle\cos(\phi_{1}-\phi_{2})\rangle/N$, which accounts qualitatively for its excess over $v_{2}\{2\}$ when multiplicity is low.
@@ IV.1 Physics Models
Figure A.2 in the appendix shows the avfd simulation results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The $\hat{q}_{2,{\rm pair}}^{2}$ is computed from the same POIs used for the $\Delta\gamma$ and $v_{2,{\rm single}}$ measurements. Three values of $n_{5}/s$ are simulated. The results are consistent with those of Ref. [36]. In avfd and all other physics models we studied, $\Delta\gamma$ appears to be linear as a function of $v_{2,{\rm single}}$. We fit the measured range of $0.20<v_{2,{\rm single}}<$ 0.2 to obtain the intercept $\Delta\gamma_{\textsc{ess}}$ at $v_{2,{\rm single}}=0$. The intercept is sensitive to CME with the $v_{2}$-induced background significantly reduced. For $n_{5}/s=0$, the intercept is consistent with zero as one would expect if the background was completely removed. For $n_{5}/s=0.1$ and 0.2, the intercepts are found to be $1.34\times 10^{-4}$ and $5.70\times 10^{-4}$, respectively. The CME signal is expected to scale with $(n_{5}/s)^{2}$. The ratio of the ESS intercept for $n_{5}/s=0.2$ to that for $n_{5}/s=0.1$ is $4.25\pm 0.13$, deviating from the expected ratio of 4 by about 2 standard deviations.
Figure A.2 shows the avfd results analyzed using the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$. The $\hat{q}_{2}^{2}$ is calculated using particles from $|\eta|<0.3$, separate from the POIs from $20.3<|\eta|<$ 2 (upper cut of 2 instead of typically 1 to increase statistics) that are used to calculate $\Delta\gamma$ and $v_{2}$. The range of dynamical variation of $v_{2}$ is significantly smaller than the range of variation in $v_{2,{\rm single}}$ of the ESS method, which is predominantly statistical. The intercept $\Delta\gamma_{\textsc{ese}}$ is obtained from a linear fit to all data points. The intercept for the $n_{5}/s=0$ case is consistent with zero, as expected. The intercepts for $n_{5}/s=0.1$ and 0.2 are $(1.63\pm 0.19)\times 10^{-4}$ and $(6.61\pm 0.19)\times 10^{-4}$, respectively. The statistical uncertainties are larger than those from ESS because of the larger range of projection to the intercept. The intercept ratio between $n_{5}/s=0.2$ and 0.1 is $4.06\pm 0.49$, consistent with the expected ratio of 4. It is interesting to note that the fit slopes from both ESS and ESE vary with $n_{5}/s$ significantly, whereas it is expected to measure the background strength. This is likely because the CME signal particles have altered the strength of background-contributing correlations in the final state [69] and thus affected the slope as well.
@@
Figure A.5 shows the ampt results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ calculated from particles within $|\eta|<0.3$ whereas POIs are taken from $10.3<|\eta|<$ 1. The intercepts are consistent with zero as expected, albeit with large statistical uncertainties. The fit errors on the slope parameter are too large to draw firm conclusions. We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities $53<|\eta|<$ 5 and $32<|\eta|<$ 3 for 200 GeV and 27 GeV collisions, respectively, keeping the same POI acceptance. The results are shown in Fig. A.5 and are found to be similar to those with midrapidity $\hat{q}_{2}^{2}$ but with larger uncertainties.
@@
Figure A.7 shows the epos4 results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ from $|\eta|<0.3$. The POI pseudorapidity acceptance is taken to be $20.3<|\eta|<$ 2 to increase statistics. The intercepts are consistent with zero as expected, albeit with large statistical uncertainties. The fit errors on the slope parameter are too large to draw firm conclusions.
@@
We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities of $53<|\eta|<$ 5 and $32<|\eta|<$ 3 for 200 GeV and 27 GeV collisions, respectively, keeping the same POI acceptance. The results are shown in Fig. A.11 corresponding to the same hydjet++ data as in Fig. A.9. The intercepts are all consistent with zero, except the 40–50% centrality bin at 27 GeV where the intercept is positive but within 2 standard deviations from zero.
@@ IV.2 Discussions on Physics Model Results
Figure 5: ESE intercepts. Shown are intercepts (upper panels) and intercepts divided by the corresponding inclusive $\Delta\gamma$ values (lower panels) from the ESE method for various physics models. For avfd ($n_{5}/s=0$), the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For ampt, epos4, and hydjet++, the results are from the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. The POI $\eta$ acceptance is $10.3<|\eta|<$ 1 for avfd, ampt, and hydjet++, and $20.3<|\eta|<$ 2 for epos4, and the $p_{\rm T}$ acceptance is $20.2<p_{\rm T}<$ 2 GeV/$c$ for all models. The $q_{2}^{2}$ is calculated using particles from $20.2<p_{\rm T}<$ 2 GeV/$c$ and $|\eta|<0.3$ for all models (and also forward/backward rapidities for ampt and hydjet++: $53<|\eta|<$ 5 at 200 GeV and $32<|\eta|<$ 3 at 27 GeV). Several points are missing as they are outside the frame but consistent with zero with large uncertainties.
Figure 6: ESS intercepts. Shown are intercepts (upper panels) and intercepts divided by the corresponding inclusive $\Delta\gamma$ values (lower panels) from the ESS method for various physics models. For avfd ($n_{5}/s=0)$, the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For ampt, epos4, and hydjet++, the results are from the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV. The POI $\eta$ acceptance is $|\eta|<1$ for avfd, ampt, and hydjet++, and $|\eta|<2$ for epos4, and the $p_{\rm T}$ acceptance is $20.2<p_{\rm T}<$ 2 GeV/$c$ for all models. The $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs.
@@
We observe in Fig. 6 that the ESS intercepts are close to zero but not always consistent with zero, although the background is suppressed by 80% or more as shown in the lower panels. Because of the statistical fluctuation nature, the $v_{2}$-induced backgrounds unnecessarily must be zero in the ESS intercept $\Delta\gamma_{\textsc{ess}}$. A model demonstration of this for the similar but simpler ES method (see Sect. II) can be found in Ref. [44]. There, the reason is clear, as elucidated in Sect. II and by the middle-panel cartoon of Fig. 1: the resonance (generally background contributing source) $v_{2}$ values are proportional to the overall $v_{2}$ of final-state particles, and when events are binned in the observed, statistically fluctuating $v_{2}^{\rm obs}$ of final-state particles, the average $\left\langle v_{2}\right\rangle$ values of the background sources are positive at $v_{2}^{\rm obs}=0$ [44]. The situation is much more complicated here in the ESS method where $\Delta\gamma$ is plotted against $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}$. As discussed in Sect. III.7, the $\hat{q}_{2,{\rm pair}}^{2}$ calculated by Eq. 9 is composed of several components: (i) a flow component from hydrodynamic particle pairs, which is related to the single particle flow in a non-trivial way [69], (ii) a flow component of the resonances (CME background contributing sources), (iii) a nonflow component between the two pairs used to calculate the two-pair cumulant by Eq. 9 that contain daughters from, for example, the same resonance decay, and (iv) a self-correlation component from the sharing of a common particle between the two pairs. Because of these complicated nature in the event-by-event $\hat{q}_{2,{\rm pair}}^{2}$ variable, it is unclear how the $v_{2,{\rm single}}$ and $\Delta\gamma$ would behave, individually and intertwiningly, as functions of $\hat{q}_{2,{\rm pair}}^{2}$. It would not be expected a priori that the $\Delta\gamma$ as a function of $v_{2,{\rm single}}$, binned in $\hat{q}_{2,{\rm pair}}^{2}$, would project to a zero intercept in the absence of CME.
@@
To conclude this part of the discussion, our physics model results show that the ESS intercept is not a well-controlled observable. Although the background in $\Delta\gamma_{\textsc{ess}}$ is strongly suppressed, one does not know how much background still remains in $\Delta\gamma_{\textsc{ess}}$, or even its sign. One does not know quantitatively what the ESS intercept exactly measures, so the ESS method is practically not useful to search for the CME.

### v4

@@ III.7 Analysis Details
Figure 3: $\hat{q}_{2}^{2}$ distributions. Shown are the single $\hat{q}_{2}^{2}$ and pair $\hat{q}_{2,{\rm pair}}^{2}$ distributions of POIs from avfd with $n_{5}/s=0$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I (lower center), and Toy Model II (lower right). Shown in all panels are for the 30–40% centrality of Au+Au collisions, as examples. The $q_{2}^{2}$ (used in the ESE analysis) is calculated from particles in $|\eta|<0.3$, and the pair $q_{2,{\rm pair}}^{2}$ (used in the ESS analysis) is calculated from particles within $|\eta|<1$, both with $0.2<p_{\rm T}<2$ GeV/$c$.
POIs are assumed to be all charged hadrons, defined to be $\pi^{\pm}$, $K^{\pm}$, pp and $\bar{p}$ in our study, and are within the transverse momentum range of $0.2<p_{\rm T}<2$ GeV/$c$ and a certain central pseudorapidity ($\eta$) range, typical of midrapidity detectors such as the STAR experiment [72].
@@
  • east subevent: $0.3<\eta<1$ (or $0.3<\eta<2$).
The POIs are taken from either the west subevent or the east subevent, but not from both. The event selection quantity $\hat{q}_{2}^{2}$ is computed by the two-particle cumulant method in Eqs. 5,6, using particles from the center subevent, $|\eta|<0.3$ (or from a forward/backward region $3<|\eta|<4$).
@@
Figure 4: Average $v_{2}^{2}\{2\}$. Shown are the two-particle cumulant $v_{2}^{2}\{2\}$ of POIs and $v_{2,{\rm pair}}^{2}\{2\}$ of particle pairs of POIs as functions of centrality from avfd for $n_{5}/s=0$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I (lower center), and Toy Model II (lower right). POIs are from acceptances $|\eta|<1$ (except for epos4, $|\eta|<2$) and $0.2<p_{\rm T}<2$ GeV/$c$. The large departure of $v_{2,{\rm pair}}^{2}\{2\}$ from $v_{2}^{2}\{2\}$, both calculated by the cumulant method (Eq. 8), is largely due to self-correlations where two pairs share a common particle. The issue is severe in peripheral collisions except for hydjet++ where the average multiplicity is significantly larger than the rest models; for reference, for 30–40% centrality of Au+Au, the average multiplicities within $|\eta|<0.5$ are 151, 169, 88, and 255 for avfd, ampt, epos4, and hydjet++, respectively.
@@
The $\hat{q}_{2}^{2}$ variable calculated from single particles has been widely used. The $\hat{q}_{2,{\rm pair}}^{2}$ variable calculated from particle pairs is relatively new [36]. As seen from Eq. 7, the average $\left\langle q_{2}^{2}\right\rangle$ is related to the two-particle cumulant $v_{2}\{2\}$. To gain insights, we calculate by Eq. 8 the $v_{2}\{2\}$ of single POIs and the $v_{2,{\rm pair}}\{2\}$ of particle pairs of POIs; In the latter case, the $\phi_{1}$ and $\phi_{2}$ in Eq. 8 are the azimuthal angles of two pairs, each formed by two POIs. The $v_{2,{\rm pair}}\{2\}$ consists of a component from pairing of “random” (hydrodynamic) particle pairs or a hydrodynamic pair with a resonance-decay pair, and a component from two-resonance $v_{2}\{2\}$. The first component reflects, in a non-trivial way, the single particle $v_{2}$ [73]. The last component is simply resonance (background contributing source) $v_{2}\{2\}$. Both $v_{2}\{2\}$ and $v_{2,{\rm pair}}\{2\}$ contain nonflow effects, for example, from resonance decays (in the case of $v_{2,{\rm pair}}\{2\}$, the two pairs both contain daughters from the same resonance decay). Figure 4 shows $v_{2}\{2\}$ (from 2 POIs) and $v_{2,{\rm pair}}\{2\}$ (from 2 pairs of POIs) as functions of centrality in different simulations of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200\text{ GeV}$. They roughly agree at high multiplicity (hydjet++, or central collisions of other models), and $v_{2,{\rm pair}}\{2\}$ becomes much larger than $v_{2}\{2\}$ at low multiplicity (peripheral collisions of models except for hydjet++). This discrepancy comes from the self-correlations where the two pairs share one common POI. To get a feeling of this effect, consider a simplified scenario with $\phi_{{\rm pair}}=(\phi_{1}+\phi_{2})/2$. For two pairs with indices $(1,3)$ and $(2,3)$ where particle 3 is shared, their correlation is $\cos 2[(\phi_{1}+\phi_{3})/2-(\phi_{2}+\phi_{3})/2]=\cos(\phi_{1}-\phi_{2})$, and the fraction of this case in the $v_{2,{\rm pair}}^{2}\{2\}$ calculation is the triplet multiplicity over the quadruplet one, which is $\sim 1/N$. Therefore, besides the common flow component, $v_{2,{\rm pair}}\{2\}$ has an extra part $\sim\langle\cos(\phi_{1}-\phi_{2})\rangle/N$, which accounts qualitatively for its excess over $v_{2}\{2\}$ when multiplicity is low.
@@ IV.1 Physics Models
Figure A.2 in the appendix shows the avfd simulation results of $\Delta\gamma$ as functions of $v_{2,{\rm single}}$ in bins of $\hat{q}_{2,{\rm pair}}^{2}$ analyzed by the ESS method in 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. The $\hat{q}_{2,{\rm pair}}^{2}$ is computed from the same POIs used for the $\Delta\gamma$ and $v_{2,{\rm single}}$ measurements. Three values of $n_{5}/s$ are simulated. The results are consistent with those of Ref. [36]. In avfd and all other physics models we studied, $\Delta\gamma$ appears to be linear as a function of $v_{2,{\rm single}}$. We fit the measured range of $0<v_{2,{\rm single}}<0.2$ to obtain the intercept $\Delta\gamma_{\textsc{ess}}$ at $v_{2,{\rm single}}=0$. The intercept is sensitive to CME with the $v_{2}$-induced background significantly reduced. For $n_{5}/s=0$, the intercept is consistent with zero as one would expect if the background was completely removed. For $n_{5}/s=0.1$ and 0.2, the intercepts are found to be $1.34\times 10^{-4}$ and $5.70\times 10^{-4}$ with negligible statistical uncertainties, respectively. The CME signal is expected to scale with $(n_{5}/s)^{2}$. The ratio of the ESS intercept for $n_{5}/s=0.2$ to that for $n_{5}/s=0.1$ is $4.25\pm 0.13$, deviating from the expected ratio of 4 by about 2 standard deviations.
Figure A.2 shows the avfd results analyzed using the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$. The $\hat{q}_{2}^{2}$ is calculated using particles from $|\eta|<0.3$, separate from the POIs from $0.3<|\eta|<2$ (upper cut of 2 instead of typically 1 to increase statistics) that are used to calculate $\Delta\gamma$ and $v_{2}$. The range of dynamical variation of $v_{2}$ is significantly smaller than the range of variation in $v_{2,{\rm single}}$ of the ESS method, which is predominantly statistical. The intercept $\Delta\gamma_{\textsc{ese}}$ is obtained from a linear fit to all data points. The intercept for the $n_{5}/s=0$ case is consistent with zero, as expected. The intercepts for $n_{5}/s=0.1$ and 0.2 are $(1.63\pm 0.19)\times 10^{-4}$ and $(6.61\pm 0.19)\times 10^{-4}$, respectively. The statistical uncertainties are larger than those from ESS because of the larger range of projection to the intercept. The intercept ratio between $n_{5}/s=0.2$ and 0.1 is $4.06\pm 0.49$, consistent with the expected ratio of 4 although the uncertainty is large. It is interesting to note that the fit slopes from both ESS and ESE vary with $n_{5}/s$ significantly, whereas it is expected to measure the background strength. This is likely because the CME signal particles have altered the strength of background-contributing correlations in the final state [73] and thus affected the slope as well.
@@
Figure A.5 shows the ampt results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ calculated from particles within $|\eta|<0.3$ whereas POIs are taken from $0.3<|\eta|<1$. The intercepts are consistent with zero as expected, albeit with large statistical uncertainties. The fit errors on the slope parameter are too large to draw firm conclusions. We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities $3<|\eta|<5$ and $2<|\eta|<3$ for 200 GeV and 27 GeV collisions, respectively, keeping the same POI acceptance. The results are shown in Fig. A.5 and are found to be similar to those with midrapidity $\hat{q}_{2}^{2}$ but with larger uncertainties.
@@
Figure A.7 shows the epos4 results analyzed by the ESE method of $\Delta\gamma$ as functions of $v_{2}$ in events classified according to $\hat{q}_{2}^{2}$ from $|\eta|<0.3$. The POI pseudorapidity acceptance is taken to be $0.3<|\eta|<2$ to increase statistics. The intercepts are consistent with zero as expected, albeit with large statistical uncertainties. The fit errors on the slope parameter are too large to draw firm conclusions.
@@
We have also used $\hat{q}_{2}^{2}$ from particles at forward/backward pseudorapidities of $3<|\eta|<5$ and $2<|\eta|<3$ for 200 GeV and 27 GeV collisions, respectively, keeping the same POI acceptance. The results are shown in Fig. A.11 corresponding to the same hydjet++ data as in Fig. A.9. The intercepts are all consistent with zero, except the 40–50% centrality bin at 27 GeV where the intercept is positive but within 2 standard deviations from zero.
@@ IV.2 Discussions on Physics Model Results
Figure 5: ESE intercepts. Shown are intercepts (upper panels) and intercepts divided by the corresponding inclusive $\Delta\gamma$ values (lower panels) from the ESE method for various physics models. For avfd ($n_{5}/s=0$), the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For ampt, epos4, and hydjet++, the results are averaged over the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV ($\left\langle\Delta\gamma_{\textsc{ese}}/\Delta\gamma\right\rangle$ does not necessarily equal $\left\langle\Delta\gamma_{\textsc{ese}}\right\rangle/\left\langle\Delta\gamma\right\rangle$). The POI $\eta$ acceptance is $0.3<|\eta|<1$ for avfd, ampt, and hydjet++, and $0.3<|\eta|<2$ for epos4, and the $p_{\rm T}$ acceptance is $0.2<p_{\rm T}<2$ GeV/$c$ for all models. The $q_{2}^{2}$ is calculated using particles from $0.2<p_{\rm T}<2$ GeV/$c$ and $|\eta|<0.3$ for all models (and also forward/backward rapidities for ampt and hydjet++: $3<|\eta|<5$ at 200 GeV and $2<|\eta|<3$ at 27 GeV). Several points are missing as they are outside the frame but consistent with zero with large uncertainties.
Figure 6: ESS intercepts. Shown are intercepts (upper panels) and intercepts divided by the corresponding inclusive $\Delta\gamma$ values (lower panels) from the ESS method for various physics models. For avfd ($n_{5}/s=0)$, the result is from the 30–40% centrality Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV. For ampt, epos4, and hydjet++, the results are averaged over the 20–50% centrality range of Au+Au collisions at $\sqrt{s_{{}_{NN}}}=200$ and 27 GeV ($\left\langle\Delta\gamma_{\textsc{ess}}/\Delta\gamma\right\rangle$ does not necessarily equal $\left\langle\Delta\gamma_{\textsc{ess}}\right\rangle/\left\langle\Delta\gamma\right\rangle$). The POI $\eta$ acceptance is $|\eta|<1$ for avfd, ampt, and hydjet++, and $|\eta|<2$ for epos4, and the $p_{\rm T}$ acceptance is $0.2<p_{\rm T}<2$ GeV/$c$ for all models. The $\hat{q}_{2,{\rm pair}}^{2}$ is always calculated from the same POIs.
@@
We observe in Fig. 6 that the ESS intercepts are close to zero but not always consistent with zero. Since these models shown in Fig. 6 contains no CME but only physics background, our results demonstrate that the $v_{2}$-induced background is unnecessarily zero in the ESS intercept $\Delta\gamma_{\textsc{ess}}$. Although the background may be suppressed, it is of no use because the amount of suppressed background is unknown, even whether it is over-suppressed or under-suppressed is unknown as shown in the lower panels of Fig. 6. It has been shown in [36] that CME signals would be preserved in the ESS intercept, however, such a conclusion or further studies of CME signal response in ESS is not helpful before the background issue is fully addressed. The reason that the $v_{2}$-induced backgrounds unnecessarily must be zero in the ESS intercept $\Delta\gamma_{\textsc{ess}}$ is because of the statistical fluctuation nature of the observables used in the ESS analysis. A model demonstration of this for the similar but simpler ES method (see Sect. II) can be found in Ref. [48]. There, the reason is clear, as elucidated in Sect. II and by the middle-panel cartoon of Fig. 1: the resonance (generally background contributing source) $v_{2}$ values are proportional to the overall $v_{2}$ of final-state particles, and when events are binned in the observed, statistically fluctuating $v_{2}^{\rm obs}$ of final-state particles, the average $\left\langle v_{2}\right\rangle$ values of the background sources are positive at $v_{2}^{\rm obs}=0$ [48]. The situation is much more complicated here in the ESS method where $\Delta\gamma$ is plotted against $v_{2,{\rm single}}$ in events binned in $\hat{q}_{2,{\rm pair}}^{2}$. As discussed in Sect. III.7, the $\hat{q}_{2,{\rm pair}}^{2}$ calculated by Eq. 9 is composed of several components: (i) a flow component from hydrodynamic particle pairs, which is related to the single particle flow in a non-trivial way [73], (ii) a flow component of the resonances (CME background contributing sources), (iii) a nonflow component between the two pairs used to calculate the two-pair cumulant by Eq. 9 that contain daughters from, for example, the same resonance decay, and (iv) a self-correlation component from the sharing of a common particle between the two pairs. Because of these complicated nature in the event-by-event $\hat{q}_{2,{\rm pair}}^{2}$ variable, it is unclear how the $v_{2,{\rm single}}$ and $\Delta\gamma$ would behave, individually and intertwiningly, as functions of $\hat{q}_{2,{\rm pair}}^{2}$. It would not be expected a priori that the $\Delta\gamma$ as a function of $v_{2,{\rm single}}$, binned in $\hat{q}_{2,{\rm pair}}^{2}$, would project to a zero intercept in the absence of CME.
@@
To conclude this part of the discussion, our physics model results show that the ESS intercept is not an easily interpretable observable. Although the background in $\Delta\gamma_{\textsc{ess}}$ is strongly suppressed, one does not know how much background still remains in $\Delta\gamma_{\textsc{ess}}$, or even its sign. One does not know quantitatively what the ESS intercept exactly measures, so the ESS method is practically not useful to search for the CME.